OpenStudy (anonymous):

What is the maximum or minimum value of the function? What is the range? y = –2x^2 + 32x –12

OpenStudy (cwrw238):

find f'(x) and equate it to 0 solve for x

OpenStudy (anonymous):

I don't understand that ?

OpenStudy (cwrw238):

oh - i've assumed you know some calculus

OpenStudy (anonymous):

Would it be like y=-2(0)^2+32(0)-12 ?

OpenStudy (cwrw238):

no what you have there is the intercept on y-axis

OpenStudy (cwrw238):

you need to change it to the vertex form y = –2x^2 + 32x –12 = -2(x^2 - 16x + 6) = -2( x - 8)^2 -64 + 6) = -2[(x-8)^2 - 58)] so the x coordinate at maximum = 8 a negative coefficient before x^3 means there is a maximum value

OpenStudy (anonymous):

oh , but it says the maximum and range is only 116 . its either 116 or -116 ?

OpenStudy (cwrw238):

at maximum the function has a value -2*-58 = 116

OpenStudy (anonymous):

ok , i see that . . .

OpenStudy (cwrw238):

the range is f(x)<= 116

OpenStudy (anonymous):

less than or greater than ?

OpenStudy (cwrw238):